# MEDTalk: Multimodal Controlled 3D Facial Animation with Dynamic Emotions by Disentangled Embedding

### v2

@@
Abstract
Audio-driven emotional 3D facial animation aims to generate synchronized lip movements and vivid facial expressions. However, most existing approaches focus on static and predefined emotion labels, limiting their diversity and naturalness. To address these challenges, we propose MEDTalk, a novel framework for fine-grained and dynamic emotional talking head generation. Our approach first disentangles content and emotion embedding spaces from motion sequences using a carefully designed cross-reconstruction process, enabling independent control over lip movements and facial expressions. Beyond conventional audio-driven lip synchronization, we integrate audio and speech text, predicting frame-wise intensity variations and dynamically adjusting static emotion features to generate realistic emotional expressions. Furthermore, to enhance control and personalization, we incorporate multimodal inputs—including text descriptions and reference expression images—to guide the generation of user-specified facial expressions. With MetaHuman as the priority, our generated results can be conveniently integrated into the industrial production pipeline.
Introduction
3D talking head generation aims to synthesize realistic facial animations from speech or text inputs and has become a significant research area in computer graphics and artificial intelligence. This technology has broad applications in virtual assistants, digital avatars, gaming, and education (Korban and Li 2022; Al Farsi et al. 2021), enhancing human-computer interaction through lifelike digital characters. Compared to traditional manual animation, which is time-consuming and labor-intensive, data-driven approaches significantly improve efficiency while producing expressive and natural facial animations.
Early studies (Edwards et al. 2016; Pham, Cheung, and Pavlovic 2017; Pham, Wang, and Pavlovic 2018; Zhou et al. 2018; Cudeiro et al. 2019; Richard et al. 2021) primarily focused on lip synchronization while neglecting movements in other facial regions, leading to unnatural animations. Recent methods have introduced emotion control by incorporating emotion or intensity labels (Fan et al. 2022; Pan et al. 2023; Liu et al. 2024), enhancing the expressiveness of facial animations. However, label-based approaches impose a fixed emotional state throughout the animation, limiting diversity and expressiveness. Some studies (Peng et al. 2023b; Kim et al. 2024) try to extract emotion features from the input audio to improve the diversity of facial expressions. However, emotional features are not fully represented in audio, reflected in inaccurate predictions. We argue that high-quality facial animation relies on two key components: lip synchronization and emotional expressiveness. Lip movements must align precisely with speech audio, while emotional expressions should exhibit both diversity and dynamic micro-expressions to enhance realism.
@@ Introduction
Related Work
Audio-driven 3D Facial Animation
@@
Early works (Pham, Wang, and Pavlovic 2018; Cudeiro et al. 2019) primarily focused on obtaining realistic, speaker-independent lip shapes from mel spectrograms using convolutional neural networks (CNNs). Transformer-based approaches, such as FaceFormer (Fan et al. 2022), leveraged wav2vec2.0 (Baevski et al. 2020) for audio feature extraction and synthesized continuous motion via autoregressive modeling, improving lip synchronization accuracy. CodeTalker (Xing et al. 2023) employed a VQ-VAE framework, treating facial animation generation as a discrete code query task, enhancing robustness. VividTalker (Zhao et al. 2023) further refined this approach by encoding head pose and mouth motion into separate discrete latent spaces, using audio features for autoregressive motion prediction. Meanwhile, SelfTalk (Peng et al. 2023a) introduced a self-supervised framework that aligns generated lip shapes with corresponding speech text, incorporating text-audio consistency as additional supervision. Recent methods (Chen et al. 2023; Stan, Haque, and Yumak 2023; Sun et al. 2024b) have explored the diffusion model (Ho, Jain, and Abbeel 2020), leveraging its generative diversity to enhance realism and motion diversity in talking head generation
@@
Emotional Talking Head
@@
Some works apply predefined emotion labels to control the generated facial expression, such as EMOTE (Daněček et al. 2023), EmoFace (Liu et al. 2024), and FlowVQTalker (Tan, Ji, and Pan 2024a). Some works have explored extracting emotion features from audio to improve expressiveness (Ji et al. 2021; Peng et al. 2023b; Tan, Ji, and Pan 2023). They applied cross-reconstruction to separate emotional cues from speech. DEEPTalk (Kim et al. 2024) constructed a joint feature space for audio and expressions, connecting the emotional features between them. AVI-Talking (Sun et al. 2024a) utilized a large language model (LLM) to analyze emotion features in speech.
Despite these advancements, most existing methods rely on static emotion features, resulting facial animations that lack variation. To address this, DEITalk (Shen et al. 2024) combined emotion labels with dynamically estimated emotion intensities derived from a speech-expression joint space. However, the predicted intensity lacked supervision and was heavily correlated with input audio. In contrast, our approach leverages both audio and textual modalities to predict dynamic emotional intensity in a supervised manner, enhancing the stability of generated emotional variations.
Multimodal Guided Generation
@@
Aligned Multi-modal Emotion encoder (Xu et al. 2023) embeds text, image, and audio in a unified space. Leveraging the rich semantic priors of CLIP, this method allows for multimodal guidance in facial expression generation. EAT (Gan et al. 2023) proposed an adaptive Emotional Deformation Network, which generates emotion-compliant talking heads based on either emotional labels or text descriptions.
Recent studies (Ma et al. 2023a; Zhong et al. 2024; Zhao et al. 2024; Tan et al. 2025b; Tan, Ji, and Pan 2024b; Tan et al. 2024, 2025a) have explored one-shot talking head generation using text descriptions or reference expression images. By extracting the talking style of multimodal guidance through a carefully designed feature encoder, or a CLIP-based style encoder (Radford et al. 2021), these methods utilized textual or visual inputs indicating desired speaking styles, enabling realistic talking head synthesis without requiring reference videos. PMMTalk (Han et al. 2024) integrates an off-the-shelf talking head generation method with speech recognition, extracting visual and textual cues from speech. It aligns audio, image, and text features at the temporal and semantic levels to guide generation. MMHead (Wu et al. 2024) extends this approach to text-guided 3D head animation, supporting text-to-3D facial motion generation by hierarchical text annotation.
Building on these advancements, our method enables text descriptions or reference images to guide facial expression synthesis, generating nuanced expressions beyond those achievable with emotion labels alone.
Methods
Overview
@@ Methods
Disentangled Embedding Space
Some prior studies (Ji et al. 2021; Peng et al. 2023b; Nocentini, Ferrari, and Berretti 2024) have attempted to separate content and emotion features from speech signals, allowing direct emotion extraction from speech for facial animation generation. However, since emotion representation in speech is inherently incomplete, even state-of-the-art speech emotion recognition (SER) models (Chen et al. 2022; Ma et al. 2023b; Tang et al. 2023; Radford et al. 2023; Chen et al. 2024; Team et al. 2024; Fu et al. 2025) fail to achieve fully accurate predictions. This uncertainty makes emotion extraction from speech unreliable for precise avatar generation applications.
@@
2.
Cross-training requires perfectly aligned sequences of equal length with the same content but different emotions. In audio, collecting such data is challenging, and Dynamic Time Warping (DTW) (Müller 2007) is required for sequence alignment, which could introduce distortions. In contrast, we can generate aligned sequences by applying identical speech content with different emotion labels.
3.
@@
To obtain data for cross-reconstruction training, we leverage EmoFace (Liu et al. 2024), a pre-trained model that takes audio and emotion labels as inputs and predicts facial animation. Specifically, we generate speech samples with seven emotions to create a large-scale dataset. It enables flexible content-emotion pairing during training.
Through our experiments, we found that mutual information minimization, such as Gradient Reversal Layer (GRL) (Ganin and Lempitsky 2015) and vCLUB (Cheng et al. 2020) used in voice conversion tasks (Yang et al. 2022; Williams et al. 2021), was insufficient for effective content-emotion separation, leading to suboptimal generation quality. In contrast, we draw inspiration from (Ge et al. 2020) and adopt a novel training scheme, as illustrated in Fig. 3, achieving high-quality feature disentanglement.
@@
Dynamic Emotional Facial Animation
@@
Prior studies (Peng et al. 2023b; Liu et al. 2024) have demonstrated that wav2vec2.0 (Baevski et al. 2020) effectively extracts lip-related features from input audio . Consequently, we extract audio content features via wav2vec2.0 and a lightweight mapping network. Our approach achieves perfect lip-sync with fast inference, making it practical for real-time applications.
During training, the predicted audio content embeddings, combined with the emotion embeddings generated by the frozen emotion encoder, are used as inputs to the frozen decoder to predict facial expressions. In addition to the reconstruction loss , we introduce a similarity loss , which measures the cosine similarity of content embedding from the audio and motion:
@@
To predict frame-wise emotional intensity, we leverage both prosodic and semantic information. Precisely, we extract frame-wise emotion-related features from audio using emotion2vec (Ma et al. 2023b), avoiding explicit emotion recognition. As for speech text, we utilize Whisper (Radford et al. 2023) to transcribe audio into text and encode using a pre-trained RoBERTa-based model (Cui et al. 2020), obtaining text embeddings . These two modalities are integrated through the Cross-Modality Fusion (CMF) module, which employs cross-attention to predict frame-wise intensity , as depicted in Fig. 2(b).
@@
Multimodal Guided Animation
@@
To guide expression generation, we collect reference images and corresponding text descriptions from a 2D talking head dataset, RAVDESS (Livingstone and Russo 2018). Specifically, Mediapipe (Lugaresi et al. 2019) is employed to extract blendshape coefficients for each frame. Then, we select the frame with the highest emotional intensity as the reference image. To obtain a text description of the facial expression, we employ Google Gemini 2.0 Flash Thinking (Google 2025), which provides a natural language interpretation of the selected frame.
@@
Given CLIP’s (Radford et al. 2021) robust text and image encoding capabilities, we freeze the CLIP text encoder and image encoder . To align these features with the emotion label embedding space, we introduce projection networks and to map the encoded features to latent features and , which have the same dimension as . In the multimodal guided generation mode, these multimodal embeddings replace the emotion label embeddings, while the rest of the model remains unchanged.
@@
Experiments
Experimental Settings
@@
Due to the limited availability of publicly accessible MetaHuman controller rig datasets, we utilize the dataset introduced by EmoFace (Liu et al. 2024) for training audio-driven expression generation. This dataset includes short-sentence utterances across seven emotional categories.
For multimodal guided animation, we employ RAVDESS (Livingstone and Russo 2018). Though EmoTalk (Peng et al. 2023b) has extracted its blendshape coefficients for each frame, they cannot be directly used for training. We utilized Unreal Engine’s Live Link Plugin to convert the blendshape coefficients into MetaHuman controller values to bridge this gap.
@@
We compare our MEDTalk with some publicly available audio-driven 3D talking-head animation models, namely FaceFormer (Fan et al. 2022), EmoTalk (Peng et al. 2023b), EmoFace (Liu et al. 2024), DiffPoseTalk (Sun et al. 2024b). As these models were trained on different datasets, we adjusted the output dimension to 174 to match our dataset and modified the style code to align with the embedding for different emotion labels. Each of the modified models was retrained on the same dataset until convergence and evaluated under identical test conditions.
Quantitative Evaluation
To evaluate the accuracy of generated animation, we propose metrics Mean Lip Error (MLE) and Mean Emotion Error (MEE) inspired by Lip Vertex Error and Emotion Vertex Error(Fan et al. 2022; Richard et al. 2021; Peng et al. 2023b). We calculate the average error between the predicted and the ground truth lip and emotion region values.
Additionally, to evaluate the dynamic characteristics of facial expressions, we introduce Emotion Intensity Error (EIE) and Upper-Face Rig Deviation (FRD). EIE computes the error between the predicted and ground truth emotion intensity per frame. FRD (Pan et al. 2025) quantifies the diversity of upper-face animations by comparing the average standard deviation between predicted and real animations. The evaluation metrics are defined as:

### v3

@@
Abstract.
Audio-driven emotional 3D facial animation aims to generate synchronized lip movements and vivid facial expressions. However, most existing approaches focus on static and predefined emotion labels, limiting their diversity and naturalness. To address these challenges, we propose MEDTalk, a novel framework for fine-grained and dynamic emotional talking head generation. Our approach first disentangles content and emotion embedding spaces from motion sequences using a carefully designed cross-reconstruction process, enabling independent control over lip movements and facial expressions. Beyond conventional audio-driven lip synchronization, we integrate audio and speech text, predicting frame-wise intensity variations and dynamically adjusting static emotion features to generate realistic emotional expressions. Furthermore, to enhance control and personalization, we incorporate multimodal inputs—including text descriptions and reference expression images—to guide the generation of user-specified facial expressions. With MetaHuman as the priority, our generated results can be conveniently integrated into the industrial production pipeline. The code is available at: https://github.com/SJTU-Lucy/MEDTalk.
1. Introduction
3D talking head generation aims to synthesize realistic facial animations from speech or text inputs and has become a significant research area in computer graphics and artificial intelligence. This technology has broad applications in virtual assistants, digital avatars, gaming, and education (Korban and Li, 2022; Al Farsi et al., 2021), enhancing human-computer interaction through lifelike digital characters. Compared to traditional manual animation, which is time-consuming and labor-intensive, data-driven approaches significantly improve efficiency while producing expressive and natural facial animations.
Early studies (Edwards et al., 2016; Pham et al., 2017, 2018; Zhou et al., 2018; Cudeiro et al., 2019; Richard et al., 2021) primarily focused on lip synchronization while neglecting movements in other facial regions, leading to unnatural animations. Recent methods have introduced emotion control by incorporating emotion or intensity labels (Fan et al., 2022; Pan et al., 2023; Liu et al., 2024), enhancing the expressiveness of facial animations. However, label-based approaches impose a fixed emotional state throughout the animation, limiting diversity and expressiveness. Some studies (Peng et al., 2023b; Kim et al., 2024) try to extract emotion features from the input audio to improve the diversity of facial expressions. However, emotional features are not fully represented in audio, reflected in inaccurate predictions. We argue that high-quality facial animation relies on two key components: lip synchronization and emotional expressiveness. Lip movements must align precisely with speech audio, while emotional expressions should exhibit both diversity and dynamic micro-expressions to enhance realism.
@@ 1. Introduction
2. Related Work
2.1. Audio-driven 3D Facial Animation
@@
Early works (Pham et al., 2018; Cudeiro et al., 2019) primarily focused on obtaining realistic, speaker-independent lip shapes from mel spectrograms using convolutional neural networks (CNNs). Transformer-based approaches, such as FaceFormer (Fan et al., 2022), leveraged wav2vec2.0 (Baevski et al., 2020) for audio feature extraction and synthesized continuous motion via autoregressive modeling, improving lip synchronization accuracy. CodeTalker (Xing et al., 2023) employed a VQ-VAE framework, treating facial animation generation as a discrete code query task, enhancing robustness. VividTalker (Zhao et al., 2023) further refined this approach by encoding head pose and mouth motion into separate discrete latent spaces, using audio features for autoregressive motion prediction. Meanwhile, SelfTalk (Peng et al., 2023a) introduced a self-supervised framework that aligns generated lip shapes with corresponding speech text, incorporating text-audio consistency as additional supervision. Recent methods (Chen et al., 2023; Stan et al., 2023; Sun et al., 2024b) have explored the diffusion model (Ho et al., 2020), leveraging its generative diversity to enhance realism and motion diversity in talking head generation
@@
2.2. Emotional Talking Head
@@
Some works apply predefined emotion labels to control the generated facial expression, such as EMOTE (Daněček et al., 2023), EmoFace (Liu et al., 2024), and FlowVQTalker (Tan et al., 2024a). Some works have explored extracting emotion features from audio to improve expressiveness (Ji et al., 2021; Peng et al., 2023b; Tan et al., 2023). They applied cross-reconstruction to separate emotional cues from speech. DEEPTalk (Kim et al., 2024) constructed a joint feature space for audio and expressions, connecting the emotional features between them. AVI-Talking (Sun et al., 2024a) utilized a large language model (LLM) to analyze emotion features in speech.
Despite these advancements, most existing methods rely on static emotion features, resulting in facial animations that lack variation. To address this, DEITalk (Shen et al., 2024) combined emotion labels with dynamically estimated emotion intensities derived from a speech-expression joint space. However, the predicted intensity lacked supervision and was heavily correlated with input audio. In contrast, our approach leverages both audio and textual modalities to predict dynamic emotional intensity in a supervised manner, enhancing the stability of generated emotional variations.
2.3. Multimodal Guided Generation
@@
Aligned Multi-modal Emotion encoder (Xu et al., 2023) embeds text, image, and audio in a unified space. Leveraging the rich semantic priors of CLIP, this method allows for multimodal guidance in facial expression generation. EAT (Gan et al., 2023) proposed an adaptive Emotional Deformation Network, which generates emotion-compliant talking heads based on either emotional labels or text descriptions.
Recent studies (Ma et al., 2023a; Zhong et al., 2024; Zhao et al., 2024; Tan et al., 2024b, c, 2025b, 2025a) have explored one-shot talking head generation using text descriptions or reference expression images. By extracting the style of multimodal guidance through a carefully designed feature encoder, these methods utilized textual or visual inputs indicating desired speaking styles, enabling realistic talking head synthesis without requiring reference videos. PMMTalk (Han et al., 2024) integrates an off-the-shelf talking head generation method with speech recognition, extracting visual and textual cues from speech. It aligns audio, image, and text features at the temporal and semantic levels to guide generation. MMHead (Wu et al., 2024) extends this approach to text-guided 3D head animation, supporting text-to-3D facial motion generation by hierarchical text annotation.
Building on these advancements, our method enables text descriptions or reference images to guide synthesis, generating nuanced expressions beyond predefined emotion labels.
3. Methods
3.1. Overview
@@
3.2. Disentangled Embedding Space
Some prior studies (Ji et al., 2021; Peng et al., 2023b; Nocentini et al., 2024) have attempted to separate content and emotion features from speech signals, allowing direct emotion extraction from speech for facial animation generation. However, since emotion representation in speech is inherently incomplete, even state-of-the-art speech emotion recognition (SER) models (Chen et al., 2022; Ma et al., 2023b; Tang et al., 2023; Radford et al., 2023; Chen et al., 2024; Team et al., 2024; Fu et al., 2025) fail to achieve fully accurate predictions. This uncertainty makes emotion extraction from speech unreliable for precise avatar generation applications.
@@ 3.2. Disentangled Embedding Space
Cross-training requires perfectly aligned sequences of equal length with the same content but different emotions. In audio, collecting such data is challenging, and Dynamic Time Warping (DTW) (Müller, 2007) is required for sequence alignment, which could introduce distortions. In contrast, we can generate aligned sequences by applying identical speech content with different emotion labels.
@@
To obtain data for cross-reconstruction training, we leverage EmoFace (Liu et al., 2024), a pre-trained model that takes audio and emotion labels as inputs and predicts facial animation. Specifically, we generate speech samples with seven emotions to create a large-scale dataset. It enables flexible content-emotion pairing during training.
Through our experiments, we found that mutual information minimization, such as Gradient Reversal Layer (GRL) (Ganin and Lempitsky, 2015) and vCLUB (Cheng et al., 2020) used in voice conversion tasks (Yang et al., 2022; Williams et al., 2021), was insufficient for effective content-emotion separation, leading to suboptimal generation quality. In contrast, we draw inspiration from (Ge et al., 2020) and adopt a novel training scheme, as illustrated in Fig. 3, achieving high-quality feature disentanglement.
@@
3.3. Dynamic Emotional Facial Animation
@@
Prior studies (Peng et al., 2023b; Liu et al., 2024) have demonstrated that wav2vec2.0 (Baevski et al., 2020) effectively extracts lip-related features from input audio . Consequently, we extract audio content features via wav2vec2.0 and a lightweight mapping network. Our approach achieves perfect lip-sync with fast inference, making it practical for real-time applications.
During training, the predicted audio content embeddings, combined with the emotion embeddings generated by the frozen emotion encoder, serve as inputs to the frozen decoder, which predicts facial expressions. In addition to the reconstruction loss , we introduce a similarity loss , which measures the cosine similarity of content embedding from the audio and motion:
@@
To predict frame-wise emotional intensity, we leverage both prosodic and semantic information. Precisely, we extract frame-wise emotion-related features from audio using emotion2vec (Ma et al., 2023b), avoiding explicit emotion recognition. As for speech text, we utilize Whisper (Radford et al., 2023) to transcribe audio into text and encode using a pre-trained RoBERTa-based model (Cui et al., 2020), obtaining text embeddings . These two modalities are integrated through the Cross-Modality Fusion (CMF) module, which employs cross-attention to predict frame-wise intensity , as depicted in Fig. 2(b).
@@
3.4. Multimodal Guided Animation
@@
To guide expression generation, we collect reference images and corresponding text descriptions from a 2D talking head dataset, RAVDESS (Livingstone and Russo, 2018). Specifically, Mediapipe (Lugaresi et al., 2019) is employed to extract blendshape coefficients for each frame. Then, we select the frame with the highest emotional intensity as the reference image. To obtain a text description of the facial expression, we employ Google Gemini 2.0 Flash Thinking (Google, 2025), which provides a natural language interpretation of the selected frame.
@@
Given CLIP’s (Radford et al., 2021) robust text and image encoding capabilities, we freeze the CLIP text encoder and image encoder . To align these features with the emotion label embedding space, we introduce projection networks and to map the encoded features to latent features and , which have the same dimension as . In the multimodal guided generation mode, these multimodal embeddings replace the emotion label embeddings, while the rest of the model remains unchanged.
@@
4. Experiments
4.1. Experimental Settings
@@
Due to the limited availability of publicly accessible MetaHuman controller rig datasets, we utilize the dataset introduced by EmoFace (Liu et al., 2024) for training audio-driven expression generation (Sec. 3.3). This dataset includes short-sentence utterances across seven emotional categories.
For multimodal guided animation (Sec. 3.4), we employ RAVDESS (Livingstone and Russo, 2018). Though EmoTalk (Peng et al., 2023b) has extracted its blendshape coefficients for each frame, they cannot be directly used for training. We utilized Unreal Engine’s Live Link Plugin to convert the blendshape coefficients into MetaHuman controller values to bridge this gap.
@@
We compare our MEDTalk with some publicly available audio-driven 3D talking-head animation models, namely FaceFormer (Fan et al., 2022), EmoTalk (Peng et al., 2023b), EmoFace (Liu et al., 2024), DiffPoseTalk (Sun et al., 2024b). As these models were trained on different datasets, we adjusted the output dimension to 174 to match our dataset. We modified the style code to align with the embedding for different emotion labels. Each of the modified models was retrained on the same dataset until convergence and evaluated under identical test conditions.
4.2. Quantitative Evaluation
To evaluate the accuracy of generated animation, we propose metrics Mean Lip Error (MLE) and Mean Emotion Error (MEE) inspired by Lip Vertex Error and Emotion Vertex Error(Fan et al., 2022; Richard et al., 2021; Peng et al., 2023b). We calculate the average error between the predicted and ground-truth lip and emotion region values.
Additionally, to evaluate the dynamic characteristics of facial expressions, we introduce Emotion Intensity Error (EIE) and Upper-Face Rig Deviation (FRD). EIE computes the error between the predicted and ground truth emotion intensity per frame (defined in Sec. 3.3). FRD (Pan et al., 2025) quantifies the diversity of upper-face animations by comparing the average standard deviation between predicted and real animations. The evaluation metrics are defined as:
@@ 6. Conclusion
7. Acknowledgement
This work is supported by National Natural Science Foundation of China (NSFC, No. 62472285 and No. 62102255), Sponsored by CCF-NetEase ThunderFire Innovation Research Funding (NO. CCF-Netease 202508 and NO. CCF-Netease 202509).